### v1

@@ 1.2. Statements of the main results
The rest of the paper is structured as follows. We start in Section 2 with basic concepts from differential geometry and Sobolev spaces on manifolds. Section 3 is devoted to proving 1.1 (Laplace–Beltrami) and 1.2 (Bochner-Laplace), while Section 4 is devoted to proving 1.3 (tangent Stokes). Section 5 concerns properties of the Stokes operator, and, finally, Section 6 contains the proofs of 1.4, 1.5 and 1.6 (tangent Navier–Stokes).
@@ 2. Preliminaries
In Section 2.1 we introduce some classical concepts from differential geometry on manifolds, in both parametric and extrinsic forms, that will be used throughout the rest of the paper. Section 2.2 is devoted to defining Sobolev spaces on manifolds and recalling some of their properties.
2.1. Differential geometry
Denote by $\boldsymbol{\nu}=(\nu_{i})_{i=1}^{d+1}$ the outward unit normal to $\Gamma$ and $\mathbf{P}:=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ the projection operator onto the tangent plane to $\Gamma$. The exterior gradient $\nabla_{M}$ acts on $C^{1}$ scalar fields $v:\Gamma\rightarrow\mathrm{R}$ and vector fields $\mathbf{v}=(v_{i})_{i=1}^{d+1}:\Gamma\rightarrow\mathrm{R}^{d+1}$ (see [67, p. 102] for the definition of $C^{m}(\Gamma)$, $m\in\mathrm{N}$), and is given by
$\nabla_{M}v=\mathbf{P}\left(\nabla v^{e}\right)\rvert_{\Gamma},\qquad\nabla_{M}\mathbf{v}=\left(\nabla\mathbf{v}^{e}\right)\rvert_{\Gamma}\mathbf{P}=\begin{bmatrix}\nabla_{M}^{T}v_{1}\\ \vdots\\ \nabla_{M}^{T}v_{d+1}\end{bmatrix}$,
where $v^{e}$ (resp. $\mathbf{v}^{e}$) denotes the constant normal extension of $v$ (resp. $\mathbf{v}$) to an $(d+1)$-dimensional tubular neighborhood $\Omega_{\delta}$ of $\Gamma$ [24, Lemma 2.8]. The exterior divergence $\operatorname{div}_{M}$ acts on $C^{1}$ vector fields $\mathbf{v}:\Gamma\rightarrow\mathrm{R}^{d+1}$ and tensor fields $\mathbf{A}:\Gamma\rightarrow\mathrm{R}^{(d+1)\times(d+1)}$, and is defined by
$\operatorname{div}_{M}\mathbf{v}=\operatorname*{tr}(\nabla_{M}\mathbf{v}),\qquad\operatorname{div}_{M}\mathbf{A}=(\operatorname{div}_{M}\mathbf{A}_{i,\cdot})_{i=1}^{d+1},$
where $\mathbf{A}_{i,\cdot}$ denotes the $i$-th row of $\mathbf{A}$. On the other hand, the covariant gradient $\nabla_{\Gamma}$ is defined by
$\nabla_{\Gamma}v=\nabla_{M}v,\qquad\nabla_{\Gamma}\mathbf{v}=\mathbf{P}\nabla_{M}\mathbf{v}=\mathbf{P}\begin{bmatrix}\nabla_{\Gamma}^{T}v_{1}\\ \vdots\\ \nabla_{\Gamma}^{T}v_{d+1}\end{bmatrix}$. (2.1)
The covariant divergence $\operatorname{div}_{\Gamma}$ acts on $C^{1}$ vector fields $\mathbf{v}:\Gamma\rightarrow\mathrm{R}^{d+1}$ and tensor fields $\mathbf{A}:\Gamma\rightarrow\mathrm{R}^{(d+1)\times(d+1)}$, and is defined by
$\operatorname{div}_{\Gamma}\mathbf{v}=\operatorname*{tr}(\nabla_{\Gamma}\mathbf{v}),\qquad\operatorname{div}_{\Gamma}\mathbf{A}=(\operatorname{div}_{\Gamma}\mathbf{A}_{i,\cdot})_{i=1}^{d+1}.$ (2.2)
It turns out that the exterior and covariant divergences coincide, i.e.
$\operatorname{div}_{M}\mathbf{v}=\operatorname{div}_{\Gamma}\mathbf{v},\qquad\operatorname{div}_{M}\mathbf{A}=\operatorname{div}_{\Gamma}\mathbf{A}.$
Definitions (2.1) and (2.2) can be equivalently stated in terms of local parametrizations of $\Gamma$. Before stating these results, we need to recall a few concepts concerning parametrizations of surfaces. We refer to [24, §2] and [10, §1.2].
Since $\Gamma$ is $C^{2}$ and compact, there exists a finite atlas $\{(\mathcal{V}_{i},\mathcal{U}_{i},\widetilde{\boldsymbol{\chi}}_{i})\}_{i=1}^{N}$, where each of the charts $\widetilde{\boldsymbol{\chi}}_{i}:\mathcal{V}_{i}\rightarrow\mathcal{U}_{i}\cap\Gamma$ are isomorphisms of class $C^{2}$ compatible with the orientation of $\Gamma$. The sets $\mathcal{V}_{i}$ are open connected subsets of $\mathrm{R}^{d}$ and the $\mathcal{U}_{i}$ are open subsets of $\mathrm{R}^{d+1}$. Without loss of generality we can assume that there exist domains $\mathcal{W}_{i}\subseteq\mathrm{R}^{d+1}$ such that $\overline{\mathcal{W}_{i}}\subseteq\mathcal{U}_{i}$ and $\{\mathcal{W}_{i}\}_{i=1}^{N}$ is still a covering of $\Gamma$ [67, Ch. 4, Lemma 10].
By using a partition of unity, in general it will be enough to consider a single chart; hence, we drop the index $i$ for ease of notation whenever no ambiguity arises. For $\mathbf{x}\in\mathcal{U}\cap\Gamma$, we write $\mathbf{y}:=\widetilde{\boldsymbol{\chi}}^{-1}(\mathbf{x})\in\mathcal{V}$. Moreover, for every (scalar, vector or tensor) function $\mathbf{f}$ defined on $\mathcal{U}\cap\Gamma$ we let $\widetilde{\mathbf{f}}:=\mathbf{f}\circ\widetilde{\boldsymbol{\chi}}$, which is defined on $\mathcal{V}$; equivalently $\mathbf{f}(\mathbf{x})=\widetilde{\mathbf{f}}(\mathbf{y})$ for each $\mathbf{y}\in\mathcal{V}$.
The first fundamental form is given by $\widetilde{\mathbf{g}}(\mathbf{y})=\nabla^{T}\widetilde{\boldsymbol{\chi}}(\mathbf{y})\nabla\widetilde{\boldsymbol{\chi}}(\mathbf{y})\in\mathrm{R}^{d\times d}$ for each $\mathbf{y}\in\mathcal{V}$. The non-degeneracy of $\widetilde{\boldsymbol{\chi}}$ on $\mathcal{V}$ ensures that $\widetilde{\mathbf{g}}=(\widetilde{g}_{ij})_{i,j=1}^{d}$ is symmetric uniformly positive definite on $\mathcal{V}$. Its inverse matrix is denote by $\widetilde{\mathbf{g}}^{-1}=(\widetilde{g}^{ij})_{i,j=1}^{d}$.
The area element $\widetilde{a}_{\boldsymbol{\chi}}$ is defined by $\widetilde{a}_{\boldsymbol{\chi}}(\mathbf{y})=\sqrt{\det\widetilde{\mathbf{g}}(\mathbf{y})}$. A measurable function $v:\mathcal{U}\cap\Gamma\rightarrow\mathrm{R}$ is integrable if and only if the measurable function $\widetilde{v}:\mathcal{V}\rightarrow\mathrm{R}$ is integrable. The area element links their integral values via the identity
@@ 2.1. Differential geometry
$\int_{\mathcal{U}\cap\Gamma}v=\int_{\mathcal{V}}\widetilde{v}\,\widetilde{a}_{\boldsymbol{\chi}}.$ (2.3)
We are now ready to state the equivalent definitions of the aforementioned projection operator and manifold differential operators in terms of the parametrization. They are given by
$\mathbf{P}(\mathbf{x})=\widetilde{\mathbf{P}}(\mathbf{y})=\nabla\widetilde{\boldsymbol{\chi}}(\mathbf{y})\widetilde{\mathbf{g}}^{-1}(\mathbf{y})\nabla^{T}\widetilde{\boldsymbol{\chi}}(\mathbf{y}),\qquad\qquad\nabla_{\Gamma}v(\mathbf{x})=\nabla\widetilde{\boldsymbol{\chi}}(\mathbf{y})\widetilde{\mathbf{g}}(\mathbf{y})^{-1}\nabla\widetilde{v}(\mathbf{y}),$ (2.4)
$\operatorname{div}_{\Gamma}\mathbf{v}(\mathbf{x})=\sum_{i,j=1}^{d}\widetilde{g}^{ij}(\mathbf{y})\partial_{i}\widetilde{\boldsymbol{\chi}}(\mathbf{y})\cdot\partial_{j}\widetilde{\mathbf{v}}(\mathbf{y}),\qquad\Delta_{\Gamma}v(\mathbf{x})=\frac{1}{\widetilde{a}_{\boldsymbol{\chi}}(\mathbf{y})}\operatorname{div}(\widetilde{a}_{\boldsymbol{\chi}}(\mathbf{y})\mathbf{g}(\mathbf{y})^{-1}\nabla\widetilde{v}(\mathbf{y})).$ (2.5)
It is sometimes convenient to write $\nabla_{\Gamma}v=(\underline{D}_{1}v,\dotsc,\underline{D}_{d+1}v)$ for the $d+1$ components of $\nabla_{\Gamma}v=\nabla_{M}v$; that is
$\underline{D}_{i}v(\mathbf{x})=\sum_{k,l=1}^{d+1}\partial_{k}\widetilde{\chi}_{i}(\mathbf{y})\widetilde{g}^{kl}(\mathbf{y})\partial_{l}\widetilde{v}(\mathbf{y}).$ (2.6)
@@
This notation allows us, in particular, to express the action of $\operatorname{div}_{\Gamma}$ and $\Delta_{\Gamma}$ in compact form:
@@
Notice that in general $\underline{D}_{i}\underline{D}_{j}v\neq\underline{D}_{j}\underline{D}_{i}v$ for each $i\neq j$. In fact, from [24, Lemma 2.6], we have that
@@
for all $i,j=1,\dotsc,d+1$; where $\mathbf{B}:=\nabla_{M}\boldsymbol{\nu}=\nabla_{\Gamma}\boldsymbol{\nu}$ is the symmetric matrix called the shape operator of $\Gamma$ (aka Weingarten map). Since $|\boldsymbol{\nu}|^{2}=1$, it follows that $\mathbf{B}\boldsymbol{\nu}=\mathbf{0}$.
Recall the identity $\mathbf{P}\operatorname{div}_{\Gamma}\nabla^{T}_{\Gamma}\mathbf{v}=\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}+\left(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2}\right)\mathbf{v}$, valid for each $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$ [40, Lemma 2.1]. We will show that, loosely speaking, said identity is still valid if we replace $\nabla_{\Gamma}^{T}\mathbf{v}$ by $\nabla_{\Gamma}\mathbf{v}$ in the left-hand side (hence obtaining the Bochner Laplacian $\Delta_{B}\mathbf{v}=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}$) as long as we test against gradients of scalar functions. This elementary observation, which relies on $\Gamma$ not having a boundary, will play a crucial role throughout this paper.
Lemma 2.1 (weak commutator identity).
Assume $\Gamma$ is of class $C^{3}$. Then for each $\phi\in C^{2}(\Gamma)$ and $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$ there holds that
$\int_{\Gamma}\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}\cdot\nabla_{\Gamma}\phi=\int_{\Gamma}\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}\cdot\nabla_{\Gamma}\phi+\int_{\Gamma}(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}\cdot\nabla_{\Gamma}\phi.$ (2.8)
Proof.
Utilizing integration-by-parts formulas [11, Lemmas B.5 and B.7], the symmetry of $\nabla_{\Gamma}\nabla_{\Gamma}u$ (cf. C.5), we first write
$\int_{\Gamma}\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}\cdot\nabla_{\Gamma}\phi=-\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}:\nabla_{\Gamma}\nabla_{\Gamma}\phi=-\int_{\Gamma}\nabla_{\Gamma}^{T}\mathbf{v}:\nabla_{\Gamma}\nabla_{\Gamma}\phi=\int_{\Gamma}\mathbf{P}\operatorname{div}_{\Gamma}\nabla^{T}_{\Gamma}\mathbf{v}\cdot\nabla_{\Gamma}\phi.$
Hence, (2.8) follows from using the aforementioned identity $\mathbf{P}\operatorname{div}_{\Gamma}\nabla^{T}_{\Gamma}\mathbf{v}=\nabla_{\Gamma}\operatorname{div}_{\Gamma}\mathbf{v}+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}$. ∎
2.2. Sobolev spaces on $\Gamma$
There are many different definitions for Sobolev spaces on manifolds (whether compact or not), either in the sense of “H” and “W” of [47], e.g. [24, Def. 2.11], [41, §3], [74, p. 2], [36, §7.1], [39, §2] and [54, p. 76]. For our purposes we find it convenient to adopt [24, Def. 2.11].
Definition 2.2 (Sobolev spaces on $\Gamma$).
Assume $\Gamma$ is of class $C^{2}$. We say that a function $u\in\mathrm{L}^{1}(\Gamma)$ has the weak extrinsic gradient $\mathbf{v}\equiv\nabla_{M}u\equiv(\underline{D}_{1}u,\dotsc,\underline{D}_{d+1}u)\in\mathbf{L}^{1}(\Gamma)$ if for every function $\varphi\in C^{1}(\Gamma)$ we have
$\int_{\Gamma}u\,\underline{D}_{i}\varphi=-\int_{\Gamma}v_{i}\,\varphi+\int_{\Gamma}\operatorname*{tr}(\mathbf{B})u\,\varphi\nu_{i},\qquad\forall i\in\{1,\dotsc,d+1\},$
For each $m\in\mathrm{N}$ and $p\in[1,\infty)$, we define the Sobolev space $\mathrm{W}^{m,p}(\Gamma)$ by
[…]
@@ 2.2. Sobolev spaces on $\Gamma$
where $\nabla_{M}^{k}$ denotes the $k$-fold weak extrinsic gradient—i.e., $\nabla_{M}^{k}u$ is the $k$-dimensional tensor defined by $(\nabla_{M}^{k}u)_{\mathbf{j}}=\underline{D}_{j_{k}}\cdots\underline{D}_{j_{1}}u$ for each $\mathbf{j}=(j_{i})_{i=1}^{k}\in\{1,\dotsc,d+1\}^{k}$. We also write $\nabla_{M}^{0}u:=u$. We endow $\mathrm{W}^{m,p}(\Gamma)$ with the norm
$\|u\|_{m,p;\Gamma}=\left(\sum_{k=0}^{m}\|\nabla_{M}^{k}u\|_{0,p;\Gamma}^{p}\right)^{1/p},$
which makes it a Banach space for $p\geq 1$ and reflexive for $p>1$; the proof follows from similar arguments from its flat domain counterpart. For $p=2$, we write $\mathrm{H}^{m}(\Gamma):=\mathrm{W}^{m,2}(\Gamma)$, which is a Hilbert space.
Notice that $\mathrm{W}^{m+1,p}(\Gamma)\subseteq\mathrm{W}^{m,p}(\Gamma)$ and $\nabla_{M}u\in\mathrm{W}^{m,p}(\Gamma)$ for each $u\in\mathrm{W}^{m+1,p}(\Gamma)$. Moreover, if $\Gamma$ is of class $C^{m}$ ($m\geq 2$), then the integration-by-parts formula [11, Lemma B.5] implies that $C^{m}(\Gamma)\subseteq\mathrm{W}^{m,p}(\Gamma)$, and for each $u\in C^{m}(\Gamma)$, its weak $m$-fold weak extrinsic gradient coincides with the usual $m$-fold extrinsic gradient that results from iterating (2.4).
We write $\mathbf{W}^{m,p}(\Gamma):=[\mathrm{W}^{m,p}(\Gamma)]^{d+1}$ and endow it with its natural product-space norm. For $p=2$, we also write $\mathbf{H}^{m}(\Gamma):=\mathbf{W}^{m,2}(\Gamma)$.
Finally, we introduce the following spaces
$\mathrm{L}^{p}_{\#}(\Gamma)$ $:=\{\textstyle q\in\mathrm{L}^{p}(\Gamma):\int_{\Gamma}q=0\},$
$\mathrm{W}^{m,p}_{\#}(\Gamma)$ $:=\mathrm{L}^{p}_{\#}(\Gamma)\cap\mathrm{W}^{m,p}(\Gamma),$

### v2

@@ 3. $\bullet$
The rest of the paper is structured as follows. We start in Section 2.1 with basic concepts from differential geometry and Sobolev spaces on manifolds. We summarize in Section 2.2 some of our results from [5] regarding the Laplace–Beltrami problem that will prove useful throughout the rest of the manuscript. Section 3 is devoted to proving 1.1 (Bochner–Laplace), while Section 4 is mostly devoted to proving 1.2 (tangent Stokes). Section 5 contains the proofs of 1.3, 1.4 and 1.5 (tangent Navier–Stokes). Finally, in Section 6 we discuss the applicability of our regularity results to other choices of Laplace operators on $\Gamma$.
@@ 2. Preliminaries
In this section, we review results from differential geometry on hypersurfaces, Sobolev spaces on $\Gamma$, prove a crucial geometric identity (cf. 2.13), and review elliptic theory results for the (LB) problem. We also reconcile several definitions of Sobolev spaces, which turn out to be useful depending on the context and regularity of $\Gamma$.
2.1. Differential geometry and Sobolev spaces on $\Gamma$
In this section we collect classical results on the calculus on manifolds (see, e.g. [23, §2], [10, §1.2] and [58]) and introduce and discuss properties of Sobolev spaces on $\Gamma$. We shall be brief, as we have discussed calculus on hypersurfaces to a great extent in [5, Section 2]. We assume connectedness for convenience, otherwise we could argue independently for each connected component. Let $\Gamma$ be a $d$-dimensional ($d\geq 2$) compact and connected manifold without boundary embedded in $\mathrm{R}^{d+1}$ of class $C^{0,1}$ (see [58, p. 102] for the definition of $C^{m}(\Gamma)$ for $m\in\mathrm{N}$, which naturally extends to $C^{m-1,1}$). Since $\Gamma$ is compact, there exists a finite atlas $\{(\mathcal{V}_{i},\mathcal{U}_{i},\widetilde{\boldsymbol{\chi}}_{i})\}_{i=1}^{N}$, where each of the charts $\widetilde{\boldsymbol{\chi}}_{i}:\mathcal{V}_{i}\rightarrow\mathcal{U}_{i}\cap\Gamma$ are isomorphisms of class $C^{0,1}$ compatible with the orientation of $\Gamma$. The sets $\mathcal{V}_{i}$ are open connected subsets of $\mathrm{R}^{d}$ and the $\mathcal{U}_{i}$ are open subsets of $\mathrm{R}^{d+1}$. An a.e. defined unit normal vector field $\boldsymbol{\nu}=(\nu_{i})_{i=1}^{d+1}:\Gamma\rightarrow\mathrm{R}^{d+1}$ to $\Gamma$ is given by $\boldsymbol{\nu}:=\mathbf{N}/|\mathbf{N}|$, where $\mathbf{N}\circ\widetilde{\boldsymbol{\chi}}:=\sum_{j=1}^{d+1}\det(\mathbf{e}_{j},\nabla\widetilde{\boldsymbol{\chi}})\mathbf{e}_{j}$ for each $i=1,\dotsc,N$; here $\{\mathbf{e}_{j}\}_{j=1}^{d+1}$ is the canonical basis of $\mathrm{R}^{d+1}$. We denote by $\mathbf{P}:=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ the projection operator onto the tangent plane to $\Gamma$.
Definition 2.1 (Sobolev spaces on $\Gamma$).
Assume $\Gamma$ of class $C^{m-1,1}$, for some $m\in\mathrm{N}$, and let $\{(\mathcal{V}_{i},\mathcal{U}_{i},\widetilde{\boldsymbol{\chi}}_{i})\}_{i=1}^{N}$ be a finite atlas of $\Gamma$. For $p\in[1,\infty)$, we define the Sobolev space $\mathrm{W}^{m,p}(\Gamma)$ by
@@ 2.1. Differential geometry and Sobolev spaces on $\Gamma$
$\mathrm{W}^{m,p}(\Gamma):=\{u:\Gamma\rightarrow\mathrm{R}\mid u\circ\widetilde{\boldsymbol{\chi}}_{i}\in\mathrm{W}^{m,p}(\mathcal{V}_{i}),\quad\forall i=1,\dotsc,N\},$
endowed with the norm
$\|u\|_{m,p;\Gamma}:=\left(\sum_{i=1}^{N}\|u\circ\widetilde{\boldsymbol{\chi}}_{i}\|_{m,p;\mathcal{V}_{i}}^{p}\right)^{1/p}.$ (2.1)
We know (e.g. [5, Proposition 3.2]) that the definition of $\mathrm{W}^{m,p}(\Gamma)$ is independent of the chosen atlas. Moreover, given two finite atlas of $\Gamma$, their induced norms (2.1) are topologically equivalent.
It is easy to see that $C^{m-1,1}(\Gamma)\subseteq\mathrm{W}^{m,p}(\Gamma)$. We write $\mathbf{W}^{m,p}(\Gamma):=[\mathrm{W}^{m,p}(\Gamma)]^{d+1}$ and endow it with its natural product-space norm. For $p=2$, we also write $\mathbf{H}^{m}(\Gamma):=\mathbf{W}^{m,2}(\Gamma)$.
For $\Gamma$ of class $C^{0,1}$ and $p\in[1,\infty)$, it is possible to define differential operators exterior gradient $\nabla_{M}$, covariant gradient $\nabla_{\Gamma}:=\mathbf{P}\nabla_{M}$, exterior divergence $\operatorname{div}_{M}$ and covariant divergence $\operatorname{div}_{\Gamma}$ (the latter two coincide) acting on elements of $\mathrm{W}^{1,p}(\Gamma)$ (and its vector- and tensor-valued versions) by extending the a.e. valid formulas for classically differentiable functions on $\Gamma$. Similarly, if $\Gamma$ is of class $C^{m-1,1}(\Gamma)$ for some $m\in\mathrm{N}$, the appropriate composition of the aforementioned operators allow us to define differential operators of order $m$ acting on elements of $\mathrm{W}^{m,p}(\Gamma)$; in particular, the second order Laplace–Beltrami operator acting on scalar functions is defined by $\Delta_{\Gamma}:=\operatorname{div}_{\Gamma}\nabla_{\Gamma}$. We omit further details and simply refer to [5] for their explicit formulas and further discussion.
@@
It is sometimes convenient to write $\nabla_{\Gamma}v=(\underline{D}_{1}v,\dotsc,\underline{D}_{d+1}v)$ for the $d+1$ components of $\nabla_{\Gamma}v=\nabla_{M}v$, where $v$ is a scalar field. This notation allows us, in particular, to express the action of $\operatorname{div}_{\Gamma}$ and $\Delta_{\Gamma}$ as
@@
For $\Gamma$ of class $C^{1,1}$, we write $\mathbf{B}:=\nabla_{M}\boldsymbol{\nu}=\nabla_{\Gamma}\boldsymbol{\nu}\in\mathbb{L}^{\infty}(\Gamma)$ called the shape operator of $\Gamma$ (aka Weingarten map). Since $|\boldsymbol{\nu}|^{2}=1$, it follows that $\mathbf{B}\boldsymbol{\nu}=\mathbf{0}$ a.e. on $\Gamma$. Notice that for $\Gamma$ of class $C^{2}$, it can be proved that $\mathbf{B}$ coincides with the Hessian matrix of the distance function to $\Gamma$, and so it is symmetric.
If $\Gamma$ is of class $C^{2}$, we can express $\nabla_{M}$ and $\nabla_{\Gamma}$ (and all other aforementioned differential operators on $\Gamma$) in a purely extrinsic—that is, independent of parametrizations—manner:
$\nabla_{\Gamma}v=\nabla_{M}v=\mathbf{P}\left(\nabla v^{e}\right)\rvert_{\Gamma},$
$\nabla_{M}\mathbf{v}=\left(\nabla\mathbf{v}^{e}\right)\rvert_{\Gamma}\mathbf{P},\qquad\nabla_{\Gamma}\mathbf{v}=\mathbf{P}\left(\nabla\mathbf{v}^{e}\right)\rvert_{\Gamma}\mathbf{P},$
where $v^{e}$ (resp. $\mathbf{v}^{e}$) denotes any extension of $v$ (resp., $\mathbf{v}$) to an $(d+1)$-dimensional tubular neighborhood $\Omega_{\delta}$ of $\Gamma$ with the following property: if $v:\Gamma\rightarrow\mathrm{R}$ is of class $C^{2}$, then so is $v^{e}:\Omega_{\delta}\rightarrow\mathrm{R}$. In this case, it holds that
$\nabla_{\Gamma}\nabla_{\Gamma}v=\mathbf{P}\nabla_{M}\nabla_{M}v=\mathbf{P}(\nabla\nabla v^{e})\rvert_{\Gamma}\mathbf{P}-((\nabla v^{e})\rvert_{\Gamma}\cdot\boldsymbol{\nu})\mathbf{B},$ (2.2)
which shows that the covariant Hessian $\nabla_{\Gamma}\nabla_{\Gamma}v$ is symmetric. An instance of such extension is given by [10, eq. (1.62)]. Notice that, if $\Gamma$ is not better than of class $C^{2}$, then the constant normal extension is not a valid choice—if $v:\Gamma\rightarrow\mathrm{R}$ is of class $C^{2}$, then $v^{e}$ is not necessarily of class $C^{2}$. Formula (2.2) seems to be well-known within the community of numerical methods for surface PDEs for $\Gamma$ of class $C^{3}$ and the constant normal extension, for which the second term in the right-hand side vanishes. We are, however, unaware of a specific reference for the proof. For completeness, we prove (2.2) in all its generality in B.2.
Notice that in general $\underline{D}_{i}\underline{D}_{j}v\neq\underline{D}_{j}\underline{D}_{i}v$ for each $i\neq j$. In fact, if $\Gamma$ and $v:\Gamma\rightarrow\mathrm{R}$ are of class $C^{2}$, we have that
@@
for all $i,j=1,\dotsc,d+1$; we refer to [23, Lemma 2.6], which states the result without a proof. For completeness, we provide a short proof in B.3.
The following result states that Sobolev spaces on $\Gamma$ behave well under multiplication by smooth-enough functions.
Lemma 2.2 (product rule on $\mathrm{W}^{m,p}(\Gamma)$).
Assume $\Gamma$ is of class $C^{m-1,1}$ for some $m\in\mathrm{N}$, and let $p\in[1,\infty)$. Then for each $\theta\in C^{m-1,1}(\Gamma)$ and $u\in\mathrm{W}^{m,p}(\Gamma)$, the product $\theta u$ belongs to $\mathrm{W}^{m,p}(\Gamma)$,
$\nabla_{M}(\theta u)=\theta\nabla_{M}u+u\nabla_{M}\theta,\quad\text{a.e.~on $\Gamma$}.$
and there exists a constant $C_{\theta}>0$, depending only on $\theta$, such that
$\|\theta u\|_{m,p;\Gamma}\leq C_{\theta}\|u\|_{m,p;\Gamma}.$
Theorem 2.3 (density).
Let $p\in[1,\infty)$ and $m\in\mathrm{N}$. If $\Gamma$ is of class $C^{m-1,1}$ (resp. $C^{m}$), then $C^{m-1,1}(\Gamma)$ (resp. $C^{m}(\Gamma)$) is dense in $\mathrm{W}^{m,p}(\Gamma)$.
It is helpful to endow $\mathrm{W}^{m,p}(\Gamma)$ with an equivalent norm that is parametrization-independent and resembles its flat-domain counterpart. This is the content of the following result.
Proposition 2.4 (equivalent norm in $\mathrm{W}^{m,p}(\Gamma)$).
Assume $\Gamma$ is of class $C^{m-1,1}$, for some $m\in\mathrm{N}$, and let $p\in[1,\infty)$. Then, the mapping
$u\mapsto\left(\sum_{k=0}^{m}\|\nabla_{M}^{k}u\|_{0,p;\Gamma}^{p}\right)^{1/p},$ (2.4)
@@
defines an equivalent norm in $\mathrm{W}^{m,p}(\Gamma)$. Here, $\nabla_{M}^{k}$ denotes the $k$-fold weak extrinsic gradient—i.e., $\nabla_{M}^{k}u$ is the $k$-dimensional tensor defined by $(\nabla_{M}^{k}u)_{\mathbf{j}}=\underline{D}_{j_{k}}\cdots\underline{D}_{j_{1}}u$ for each $\mathbf{j}=(j_{i})_{i=1}^{k}\in\{1,\dotsc,d+1\}^{k}$; we also write $\nabla_{M}^{0}u:=u$.
Throughout the rest of the paper, it will be usually convenient to utilize (2.4) as the “main” norm on $\mathrm{W}^{m,p}(\Gamma)$ and consequently we will also denote it by $\|\cdot\|_{m,p;\Gamma}$ when no confusion arises. Notice that this equivalent norm naturally induces a notion of Sobolev norms on subsets of $\Gamma$: let $\gamma\subseteq\Gamma$ be a measurable set with respect to the Lebesgue measure on $\Gamma$, we write
$\|v\|_{m,p;\gamma}:=\left(\sum_{k=0}^{m}\|1_{\gamma}\,\nabla_{M}^{k}v\|_{0,p;\Gamma}^{p}\right)^{1/p},\qquad\forall v\in\mathrm{W}^{m,p}(\Gamma),$
where $1_{\gamma}$ denotes the characteristic function of $\Gamma$.
We now compare adopted definition 2.1 of Sobolev spaces and a few others found in the literature.
Remark 2.5 ($\mathrm{W}^{m,p}(\Gamma)$ coincides with $H^{p}_{m}(\Gamma)$ of Hebey and Robert [35]).
2.3 and 2.4 shows that $\mathrm{W}^{m,p}(\Gamma)$ (cf. 2.1) coincides with the space $H^{p}_{m}(\Gamma)$ of [35, Definition 2.1, Proposition 2,2].
Proposition 2.6 (integration by parts on $C^{1,1}$-manifolds).
If $\Gamma$ is of class $C^{1,1}$, then for each $u\in\mathrm{W}^{1,1}(\Gamma)$ and $\boldsymbol{\varphi}\in\mathbf{C}^{1}(\Gamma)$ it holds that
$\int_{\Gamma}u\,\operatorname{div}_{\Gamma}\boldsymbol{\varphi}=-\int_{\Gamma}\nabla_{\Gamma}u\cdot\boldsymbol{\varphi}+\int_{\Gamma}\operatorname*{tr}(\mathbf{B})u\boldsymbol{\varphi}\cdot\boldsymbol{\nu}.$ (2.5)
This is consistent with the definition of Sobolev spaces of Dziuk and Elliott [23] for $\Gamma$ of class $C^{2}$. We proved (2.5) for $\Gamma$ of class $C^{1,1}$ in [5, Proposition 3.9].
We introduce the following spaces
$\mathrm{W}^{m,p}_{\#}(\Gamma)$ $:=\{\textstyle q\in\mathrm{W}^{m,p}(\Gamma):\int_{\Gamma}q=0\},$